OpenStudy (anonymous):

It takes Nim 2 2/3 hours to weave a basket. He worked Monday through Friday, 8 hours a day. How many baskets did he make?

OpenStudy (anonymous):

divide

OpenStudy (anonymous):

Yes I did but I don't get a whole number

OpenStudy (anonymous):

\[40\div 2\tfrac{2}{3}\]

OpenStudy (anonymous):

round down then

OpenStudy (anonymous):

hmm but i think you do get a whole number

OpenStudy (anonymous):

But it asks how many baskets did he make.

OpenStudy (anonymous):

8 / 2.2/3 = 3 Monday-Friday= 5 days 3*5= 15 baskets

OpenStudy (anonymous):

\[40\div 2\tfrac{2}{3}=40\div \frac{8}{3}\]\[=40\times \frac{3}{8}\] is a whole number

OpenStudy (anonymous):

\[=40\times \frac{3}{8}=\cancel{40}^5\times \frac{3}{\cancel{8}^1}=5\times 3\]

OpenStudy (anonymous):

Did he make 15?

OpenStudy (anonymous):

yes `